# Semantic Latent Motion for Portrait Video Generation

### v1

@@ 4.3 Ablation Study
Figure 7: Unmasked Motion AutoEncoder training retains details in latent motion, while masking reduces them.
Mask Strategy. In Fig. 10 We compared the impact of using a mask strategy during the training of both the Motion AutoEncoder and Motion Generator. In training the Motion AutoEncoder, we applied a mask to both the reference and target image features, sampling the mask ratio from 0% to 90%. While reconstruction metrics showed similar results with or without the mask, we found that the version without the mask included more detailed motion information, aiding the diffusion decoder. However, the mask did not significantly affect reconstruction outcomes, and even with the mask, good results were achieved. When replacing the reference image with a different person, the version without the mask captured irrelevant details like background and clothing, while the masked version focused more on motion-related details from the reference image, aligning with our goal of abstracting motion while maintaining reconstruction quality. In training the Motion Generator, we applied a mask of 0-20% to reduce the generator’s reliance on detailed information from the previous motion and encourage it to extract details from the reference motion. Without the mask, the model struggled to generalize to other datasets, but with the mask, its generation capabilities improved and it became more adaptable to new datasets.
Table 7: Ablation studies of the Mask Strategy
Method HDTF
[…]
w/ mask 21.431 136.738 0.043 29.120 0.871
w/o mask 21.326 136.229 0.041 29.061 0.876
Figure 8: User study. The numbers in the figure indicate the win rate for each dimension.
Figure 9: Unmasked motion encoding causes overfitting to details, compromising cross-dataset generalization.
5 Discussion and Limitation
Figure 10: Motion encoding in SeMo requires minimal information, achieving good results even when 90% of the data is masked.
Figure 11: SeMo can be adapted to general video scenes.
Motion encoding requires minimal information. As demonstrated in the MAE [14], model needs very minimal to generate meaningful visual content. However, in the context of motion encoding, even less information is required. The core difference between motion encoding and image encoding lies in the fact that motion encoding has a reference image. This makes much of the detailed information in the target image unnecessary. Previous methods did not make good use of the information in the reference image. In the example Fig. 10, we observe that even 90% of the image is masked, the original video can still be reconstructed.
Transfer to General Scenarios. We investigated whether Motion AutoEncoder is applicable to general videos. In Fig. 11, We found that in single-subject motion scenarios, Motion AutoEncoder performs well in reconstruction, even effectively learning hand movements. However, in complex multi-subject motion scenarios, noticeable distortions occur. We reproduced IMF and train them on WebVid.
Abstract Resoning and Detail Generation. Current video generation models, like Sora, perform temporal reasoning and pixel generation simultaneously, neglecting redundancy in motion representation. Consider how humans think – we don’t fully visualize the next state, but rather abstract conceptual representations for efficient reasoning. Video generation models could adopt similar principles: by constructing a more compact informational space through abstraction, we enable efficient inference. Subsequently, the abstract information guides the generative model in pixel synthesis, making prediction more efficient and manageable.
Limitation. We found that the generated results have blurry teeth, and high-frequency areas like beards tend to flicker. We believe this is caused by the factor that: we resize data to 256x256 and compression into the latent space via SD VAE, there was significant information loss. However, our method is scalable, and we can easily extend it to higher resolutions, by maintaining the current approach.
6 Conclusion
In this work, we have proposed a fully self-supervised framework for portrait video generation, structured around three key steps: Abstraction, Reasoning, and Generation. First, we employ a masking strategy and a video reconstruction task to compress motion into a compact, abstract space (1D token). Second, within this space, we perform long-term modeling and efficient reasoning. Finally, the inferred motion serves as a conditioning signal to guide the diffusion model in generating fine-grained details. Furthermore, our findings indicate that motion can be further compressed, enabling a more compact and efficient abstract representation. Beyond portrait video generation, our pipeline is broadly applicable to general video generation tasks. In future research, we aim to explore its feasibility across a wider range of scenarios.

### v2

@@ 4.3 Ablation Study
Dimension and Token Number. In table 3, we compared the impact of latent motion token dimensions and token numbers on model performance. For the token dimension, we fixed the number at 1 and varied its size, finding that even with a dimension as small as 32, the model could still generate reasonable images. However, smaller dimensions resulted in reduced detailed information, requiring the motion decoder to compensate, which introduced some randomness in the output, though the layout remained consistent. For the token number, we fixed the motion representation space at 512 floats and found that increasing the number of tokens improved performance. This is due to the increased parameters used by the model to analyze motion, enabling more accurate mapping from semantic tokens to low-level pixels.
Figure 7: (a) Training loss of Stage 2 with and without mask. (b) UMap [22] visualization of latent motion. (c) Attention Map between latent motion and reference image.
Masking. In  fig. 7(a), we find that masked training accelerates Stage 2 convergence, lowers loss, and improves output quality. We infer that masking forces the model to prioritize high-level motion semantics over local details, making generation more predictable. In  fig. 7(b), we observe that masking encourages discrete latent motion features, aligning with findings in [46] where discrete distributions enhance generation. In  fig. 7(d), We observe that, guided by the latent motion, the model mainly focuses on regions such as the eyes and mouth, where the movements are more significant. In  fig. 9(c), we observe that using a masking strategy can partially separate identity and motion: replacing the reference video with another identity shows less identity leakage when masking is applied. We attribute this effect to the randomly sampled mask ratios, which encourage the model to abstract at high ratios and retain some facial details at lower ones.
5 Conclusion and Limitations
Portrait video generation has a wide range of applications, but it faces challenges in both generation efficiency and quality. Moreover, existing methods often rely on prior knowledge, which limits their applicability to fixed facial scenarios. In this work, we address these issues by proposing a self-supervised motion representation framework SeMo, which is clear and straightforward, consisting of three steps: Abstraction, Reasoning, and Generation. Thanks to the compact and expressive nature of SeMo, we demonstrate that SeMo significantly improves generation efficiency and quality. Beyond portrait video generation, our pipeline is broadly applicable to general video generation tasks. In future work, we plan to further explore its feasibility and effectiveness across a wider range of scenarios.
Limitation. Due to the need for efficient generation and better user accessibility, we limited our experiments to a resolution of 256×256. However, this downsampling, together with compression by the SD VAE, causes some information loss, resulting in blurry teeth and flickering in high-frequency regions such as beards. In future work, we plan to explore higher resolutions to further improve the visual quality of the generated results.
Appendix A Motivation
Our motivation stems from the observation that video reasoning is inherently high-dimensional and complex; predicting every pixel is highly uncertain, and humans do not explicitly forecast each pixel in their brains. Instead, we tend to abstract concepts from what we see and perform efficient reasoning in this conceptual space, which not only improves the reliability but also the efficiency of predictions. Inspired by this, we apply a similar principle to portrait video generation. Rather than directly predicting pixels, the model first abstracts each frame into high-level concepts, enabling more effective prediction within this compact space. Leveraging the strong generative and conditional capabilities of diffusion models, we then progressively refine predictions from high-level semantic concepts to final pixel-level outputs in a coarse-to-fine manner.
As illustrated in the 8, given a reference image, reference motion, and target motion, the model can easily generate the target frame at the pixel level. To this end, we combine self-supervised learning and diffusion-based generative modeling in a three-step framework: Abstraction, reasoning, and generation.
Figure 8: Motivation.
Appendix B Implementation Details
As described in Sec. 3.1, SeMo adopts an EncoderDecoder architecture. Specifically, In the motion encoder, we employ 8 Transformer layers. The motion decoder consists of 12 denoising layers, including both Motion Control Layers and Temporal Align Layers. For the motion generator, we use 8 DiT layers. Throughout all modules, the hidden dimension is set to 768 and we use 8 attention heads. The Motion Autoencoder comprises 450 million parameters, and the motion generator comprises 200 million parameters. For input preprocessing, all images and video frames are resized to $256\times 256$ resolution and pixel values are normalized to the [-1,1] range. All models are trained under similar settings: a base learning rate of $10^{-4}$ per 64 batch size and the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\text{decay}=0.05$. Learning rate is set to $1e{-4}$. In Stage1, Models are trained with a batch size of 8 on 8 NVIDIA A800 GPUs for 300K iterations. In Stage2, Models are trained with a batch size of 8 on 8 NVIDIA A800 GPUs for 200K iterations. We employ gradient clipping with a threshold of 1.0 to stabilize training.
Table 4: Training Configuration
Parameter Value
Input Video Resolution 256
Input Video Frames 16
Input Video FPS 25
Optimizer Adam; $\beta_{1}=0.9,\beta_{2}=0.95$
Learning Rate $1\times 10^{-4}$
Warmup Steps 5000
Learning Rate Scheduler Cosine Annealing
$\mathcal{L}_{KL}$ 0.001
Weight Decay 0.0001
Training Batch Size 8
Training Device 8 $\times$ 80G A800 GPUs
Appendix C Userstudy
We randomly selected 10 images from HDTF and DH-FaceVid-1K, and generated 20 videos using 2 random audio clips. We invited 50 participants, divided the participants into 5 groups, each participant was asked to select the best model from 4 different perspectives (realism, lip-sync, smooth, overall). We significantly outperform competitors in realism and match the state-of-the-art (SOTA) in lip sync, smoothness, and overall quality.
Appendix D Additional Analysis
D.1 Analysis on Mask
Additional results for the cross-ID driving task are shown in fig. 9, demonstrating that the mask strategy significantly suppresses identity leakage, although complete disentanglement is still not achieved. However, we emphasize that such disentanglement is not strictly necessary for video generation tasks. Excessive retention of fine-grained details may hinder the generation process in the second stage, while discarding unnecessary details helps stabilize both training and inference.
As shown in  fig. 10, even when 90% of the content is masked, the model is still able to reconstruct reasonable frames. This indicates a high degree of redundancy in motion information, suggesting that effective motion modeling can be achieved with only a small fraction of inputs.
Furthermore,  fig. 11 reveals that without the masking strategy, the model tends to memorize excessive details, which compromises its ability to generalize to out-of-distribution data. This observation further highlights the importance of abstract representation for robust video generation.
D.2 More Visualization Result of Attention Map
In fig. 14, We visualized the attention map between the motion latent and the reference image in the decoder. The results show that the interaction between the motion latent and the reference image mainly focuses on facial regions, especially those areas that typically undergo significant movement. In addition to the face, the model also attends to local regions prone to motion, such as earrings. By contrast, the model pays less attention to the background and other regions that are unlikely to move. This phenomenon indicates that the motion latent contains rich and detailed motion information, and the model can effectively capture and utilize these highly motion-relevant local features, thereby enabling precise motion generation in target areas.
D.3 More Visualization Result of Audio-driven Video Generation
We provide more experimental results of talking head in fig. 12 fig. 13. In  fig. 13, we observe that in extreme speech cases, such as silence, the labial sound ’emmm’ or certain interjections, other methods remain static, whereas our method exhibits natural behaviors, such as blinking during silence and producing natural mouth movements during interjections.
D.4 Transfer to general video reconstruction tasks
Since our framework is fully self-supervised, in fig. 15, we further explored its applicability to general video scenarios. We found that our method is suitable for general video scenes with clear subjects and relatively simple motion. However, in scenarios with multiple or complex subjects, trailing artifacts may occur, as shown in the figure. Compared to similar methods, our approach achieves better performance on general video datasets. In our experiments, we reproduced the IMF method and trained it on the WebVid dataset.
Appendix E Novelty
The work most closely related to ours is IMF [9]. However, our objectives and implementation strategies are fundamentally different. First, regarding objectives, IMF focuses solely on image-to-image transfer, whereas our goal is to transfer from images to videos. Moreover, the way we abstract and represent motion also differs: we employ a transformer architecture along with a masking strategy, enabling the model to encode compact and semantically rich motion features. Second, our decoding approach is distinct. We use rectified flow model as the decoder, which leads to a simpler loss design and achieves better visual results. Finally, we demonstrate the effectiveness of our method on the challenging downstream task of audio-driven video generation.
Figure 9: Cross Identity Driving Example
Figure 10: Reconstruction with Different Mask Ratio.
Figure 11: Result of the analysis on generalization.The mask strategy enhances the model’s generalization ability.
Figure 12: Visualization of Audio-driven Portrait Video Generation.
Figure 13: Visualization of Audio-driven Portrait Video Generation. It can also produce natural movements—such as blinking—even during interjections or periods of silence
Figure 14: Visualization of Attention Map.
Figure 15: Transfer to general video reconstruction tasks.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: This paper proposes a self-supervised learning based framework for portrait motion representation, which improves the efficiency and quality of portrait video generation. The proposed algorithm are illustrated in Section 3. Experimental results are illustrated in Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We include the limitations of our work in Section 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is not a theoretical paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Datasets, models, and hyperparameters used in implementing proposed algorithms are all described in detail. See Section 4 and Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We will provide the code to train and evaluate the proposed algorithm, which reproduces the experiment results in the paper after the rebuttal process.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the details of our experiments in Section 4 and Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report standard error.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the compute resources during training in Section 4 and Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We followed the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This work focuses on a academic, publicly-available benchmark. This work is not related to any private or personal data, and there’s no explicit negative social impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: our work poses no risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: This paper cite the original papers such as dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: This paper will release code for running experiments and it is well documented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM is used only for writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.